Walker & Campsey wants to invest in a new computer system, and management has narrowed the choice to Systems A and B.

System A requires an up-front cost of $125,000, after which it generates positive after-tax cash flows of $80,000 at the end of each of the next 2 years. The system could be replaced every 2 years, and the cash inflows and outflows would remain the same.

System B also requires an up-front cost of $125,000, after which it would generate positive after-tax cash flows of $60,000 at the end of each of the next 3 years. System B can be replaced every 3 years, but each time the system is replaced, both the cash outflows and cash inflows would increase by 5%.

The company needs a computer system for 6 years, after which the current owners plan to retire and liquidate the firm. The company's cost of capital is 12%. What is the NPV (on a 6-year extended basis) of the system that adds the most value?

Using the information from the problem on Walker & Campsey, what is the equivalent annual annuity (EAA) for System B? Enter your answer rounded to two decimal places.
Answer #1
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=Cost of Capital=12% 0.12
N=Year of Cash Flow
CASH FLOWS OF SYSTEM    A
N Year 0 1 2 3 4 5 6
A Cashflows for the Initial investment ($125,000) $80,000 $80,000
B Cashflows for the First Replacement ($125,000) $80,000 $80,000
C Cashflows for the Second Replacement ($125,000) $80,000 $80,000
D=A+B+C Net Cash Flow ($125,000) $80,000 ($45,000) $80,000 ($45,000) $80,000 $80,000 SUM
PV=D/(1.12^N) Present Value (PV) of Cash Flow: ($125,000) $71,429 -$35,874 $56,942 -$28,598 $45,394 $40,530 $24,824
NPV=sumof PVs Net Present Value(NPV) of System A $24,824
CASH FLOWS OF SYSTEM    B
Cash Flows of first Replacement:
Outflow=125000*1.05 $131,250
Inflows =60000*1.05 $63,000
N Year 0 1 2 3 4 5 6
A Cashflows for the Initial   investment ($125,000) $60,000 $60,000 $60,000
B Cashflows for the Replacement ($131,250) $63,000 $63,000 $63,000
D=A+B Net Cash Flow ($125,000) $60,000 $60,000 ($71,250) $63,000 $63,000 $63,000 SUM
PV=D/(1.12^N) Present Value (PV) of Cash Flow: ($125,000) $53,571 $47,832 -$50,714 $40,038 $35,748 $31,918 $33,392
NPV=sumof PVs Net Present Value(NPV) of System B $33,392
NPV of the Systemthat adds most value $33,392 (SYSTEM B)
Equivalent Annual Annuity (EAA) $8,121.79 Using PMT function of excel with Rate=12%, Nper=6,Pv=-33392)
Excel Command:PMT(12%,6,-33392)